If a 300 kg solution of 3% sugar concentration evaporates at the rate of 5 kg per hour, how many minutes will it take for the solution to become a 5% sugar-concentrated solution?
(A) 12
(B) 24
(C) 624
(D) 1240
(E) 1440

Answer by ikleyn(52832)

.
If a 300 kg solution of 3% sugar concentration evaporates at the rate of 5 kg per hour,
how many minutes will it take for the solution to become a 5% sugar-concentrated solution?
(A) 12
(B) 24
(C) 624
(D) 1240
(E) 1440
~~~~~~~~~~~~~~~~~~~~

In this problem, concentration is the ratio of weight of the sugar to the weight of the mixture.


So, initial mixture contains  0.03*300 = 3*3 = 9 kg of sugar.


We want it would be  9/M = 5% = =.05 after evaporation, where M
is the final mass of the mixture  M = 9/0.05 = 900/5 = 180 kilograms.


So, 300-180 = 120 kilograms of water should evaporate.


At the rate of 5 kg per hour, it will take  120/5 = 24 hours = 24*60 = 1440 minutes.    ANSWER

Solved and explained.

--------------

Hope, this my solution will give a good charge to your brain.